40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 26     Entry time: Mon Oct 29 12:20:15 2007
Author: waldman 
Type: Configuration 
Category: OMC 
Subject: Changed OMS filters 
I changed the OMS configuration so that some of the OMC-SUS LED channels go to a breakout box so that we can input the PDH error signal. After lunch, we will try to lock the cavity with a PDH error signal and digital filters. Then its on to dither locked stuff. Note that this LED business will have to be changed back some day. For now, it should be extremely visible because there are dangling cables and a hack job interface lying around.
ELOG V3.1.3-